# Complements of discriminants of real parabolic function singularities. II

### v1

@@ 4. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at the $X_{9}$ and $J_{10}$ singularities proves all of the conjectures in [11] concerning the number of local Petrovskii lacunas at the corresponding singular points of wavefronts.
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity when the dimension $N$ and the index $i_{+}$ are both even is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The four lacunas at the $X_{9}^{2}$ singularity, with even $N$ and odd $i_{+}$, are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (18) to these two. The unique lacuna at the $J_{10}^{3}$ singularity, with even $N$ and odd $i_{+}$, is shown in picture $(m)$ of Fig. 5.

### v2

@@ 5. Petrovskii lacunas
The above enumeration of the connected components of the discriminant complements at parabolic singularities provides a complete list of local Petrovskii lacunas at the corresponding singular points of wavefronts of hyperbolic PDEs. In this theory, generating families of functions play a crucial role. Wave functions are defined on the parameter spaces of these families via integrals along certain cycles in the zero-level sets of the corresponding functions. Wavefronts are the discriminant sets of these families. Outside of the wavefront, the integration cycles and integrand forms depend regularly on the parameters, and the wave function is regular as well. However, over the discriminant sets of parameters, the level sets containing the integration cycles degenerate and the wave function may be singular. The singularity of a wave function in a neighborhood of a wavefront point is characterized by the asymptotics of the wave function as the non-discriminant parameter values approach the point. The type of this asymptotic behavior depends primarily on the singularity class of the corresponding discriminant function and on the homology class of the integration cycle. The homology classes depend on the wavefront complement components from which we approach the wavefront point. Such a component is called a local lacuna if the wave function can be continued from it to a regular function on entire neighborhood of the wavefront point.
If the singularity of the discriminant function corresponding to a wavefront point is isolated, then whether a neighboring component of the discriminant complement is a local lacuna is determined by the topological data contained in the virtual function associated with an arbitrary function from that component (see, for example, [16]).
Versal deformations of stably equivalent function singularities can be naturally identified with each other. This identification establishes a one-to-one correspondence between the components of their discriminant complements. The set of local lacunas at a discriminant point depends on the stable equivalence class of the corresponding singularity, on the parity of the number of variables, and on the parities of the inertia indices of the quadratic part of the singularity. For example, the functions $f(x_{1},\dots,x_{n})$ and $\tilde{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}+x_{n+2}^{2}$ have the same set of local lacunas. On contrary, the function $\check{f}(x_{1},\dots,x_{n+2})\equiv f(x_{1},\dots,x_{n})+x_{n+1}^{2}-x_{n+2}^{2}$ or the stably equivalent functions in n+2k1n+2k-1 variables generally has a different set.
Theorem 46.
The numbers of local lacunas near parabolic singularities of wavefronts in versal generating families are given in Table 4.
Table 4. Numbers of local lacunas at parabolic singularities
Singularity NN even NN even NN odd NN odd
class $i_{+}$ even $i_{+}$ odd $i_{+}$ even $i_{+}$ odd
$P_{8}^{1}$ 0 0 22 0
$P_{8}^{2}$ 0 0 33 0
$X_{9}^{+}$ 11 0 22 0
$X_{9}^{-}$ 11 0 0 22
$X_{9}^{1}$ 0 0 0 0
$X_{9}^{2}$ 0 44 0 0
$J_{10}^{3}$ 0 11 0 0
$J_{10}^{1}$ 0 0 0 0
In comparison with the results of [17], this table replaces the lower estimate $\geq 2$ for $P_{8}^{2}$ with the exact number 3, and replaces some five estimates of the form $\geq k$ with the same numbers $k$.
In our current terms, the unique lacuna of the $X_{9}^{+}$ singularity for even $N$ and $i_{+}$ is component $(a)$ of Fig. 1. The two lacunas of the $X_{9}^{+}$ singularity for odd $N$ and even $i_{+}$ are the components $(a)$ and $(b)$ of the same figure. The four lacunas at the $X_{9}^{2}$ singularity with even $N$ and odd $i_{+}$ are shown in picture $(i)$ of Fig. 3, or in its rotation by the angle $\frac{\pi}{2}$, or in the pictures obtained by applying the operation (17) to these two. The unique lacuna at the $J_{10}^{3}$ singularity with even $N$ and odd $i_{+}$ is shown in picture $(m)$ of Fig. 5. The two local lacunas of the $P_{8}^{1}$ singularity at odd $N$ and even $i_{+}$ are the components with $\mbox{Ind}=1$. Three local lacunas of the $P_{8}^{2}$ singularity with odd $N$ and even $i_{+}$ are also the components with $\mbox{Ind}=1$. The only one new local lacuna found now is the component with $\mbox{Card}=262$ for $P_{8}^{2}$ singularity.